QUESTION 23

  1. A share of common stock just paid a dividend (D0) of $1.50. If the expected long-run growth rate for this stock is 5%, and if investors' required rate of return is 11.5%, what is the current stock price?

    1.

    $17.57

    2.

    $24.23

    3.

    $17.13

    4.

    $18.01

    5.

    $16.28
Answer #1
2.

$24.23

As per dividend discount model,
Current share price = D0*(1+g)/(Ke-g) Where,
= 1.50*(1+0.05)/(0.1150-0.05) D0 = $       1.50
= $    24.23 g = 5%
Ke = 11.5%